OpenStudy (anonymous):

I have a simple proof question I can not figure out. Prove: If a/b < c/d, then a/b < (a+c)/(b+d) < c/d not sure where to start. :s thanks for any help

OpenStudy (freckles):

so there are no restrictions given on a,b,c, and d?

hartnn (hartnn):

a/b < c/d ad < bc now add ab on both sides, what do u get?

hartnn (hartnn):

ofcourse, assuming all a,b,c,d are non-zero

hartnn (hartnn):

then ad < bc add cd on both sides. combine the 2 results and you'll be done :)

OpenStudy (anonymous):

Alright, simple enough. Let's make the starting equation a little simpler. \[\frac{ a }{ b } < \frac{c}{d}\] \[ad<bc \] So that's Inequality #1. For now. Let's move on to Inequality #2, which is a bit more ambiguous. You can either go left, or right. It's better if you go both ways and prove both sides. Let's go left first. \[\frac{ a }{ b } < \frac{a+c}{b+d}\] Simplifying gives us \[ab+ad<ab+cb\] Cancel them ab's out. And voila! \[ad<bc\] Compare that to Inequality #1. It's the same, right? You've proven it. But not fully. Let's take \[\frac{ a+c }{ b+d } < \frac{ c }{ d }\] Cross multiply them, you get \[ad+cd<cb+cd\] The cd's cancel out, giving you \[ad<bc\] Compare this to Inequality #1 we obtained earlier. Pretty much the same, right? Mission accomplished. Apollo, we are go for launch! Hope this helped. -swas

OpenStudy (anonymous):

Thanks so much for all of your help! :)

OpenStudy (anonymous):

Anytime.